# ChildMandarin: A Comprehensive Mandarin Speech Dataset for Young Children Aged 3-5

### v2

@@ Abstract
Automatic speech recognition (ASR) systems have advanced significantly with models like Whisper, Conformer, and self-supervised frameworks such as Wav2vec 2.0 and HuBERT. However, developing robust ASR models for young children’s speech remains challenging due to differences in pronunciation, tone, and pace compared to adult speech. In this paper, we introduce a new Mandarin speech dataset focused on children aged 3 to 5, addressing the scarcity of resources in this area. The dataset comprises 41.25 hours of speech with carefully crafted manual transcriptions, collected from 397 speakers across various provinces in China, with balanced gender representation. We provide a comprehensive analysis of speaker demographics, speech duration distribution and geographic coverage. Additionally, we evaluate ASR performance on models trained from scratch, such as Conformer, as well as fine-tuned pre-trained models like HuBERT and Whisper, where fine-tuning demonstrates significant performance improvements. Furthermore, we assess speaker verification (SV) on our dataset, showing that, despite the challenges posed by the unique vocal characteristics of young children, the dataset effectively supports both ASR and SV tasks. This dataset is a valuable contribution to Mandarin child speech research and holds potential for applications in educational technology and child-computer interaction. It will be open-source and freely available for all academic purposes.
@@ 2.1 Child Speech Recognition Corpora in Mandarin Chinese
Gao et al. (Gao et al., 2012) collected the CASS CHILD dataset, which contains 631 hours of speech from 23 children aged 1 to 4 years. However, only about 80 hours of this dataset are labeled with transcriptions, and, critically, the dataset is not publicly accessible. This restricts its use in ASR experiments and highlights the difficulty of obtaining child speech corpora in Mandarin.
@@ 3.1 Dataset details
The dataset consists of 41.25 hours of speech data with carefully crafted manual transcriptions, collected from Mandarin-speaking children aged 3 to 5 years. The gender distribution is balanced across all age groups. To ensure geographic coverage, speakers were selected from different regions of China, excluding dialectal speech. A total of 397 speakers participated, representing 22 out of 34 provincial-level administrative divisions. Accents were classified into three categories: heavy (H), moderate (M), and light (L).
@@
The content of the speech recordings was unrestricted, focusing on age-appropriate daily communication while excluding sensitive topics like violence, politics, or privacy. Manual annotations were performed by professional transcribers, who meticulously adhered to the audio content, including stutters, disfluencies, and developmental speech patterns. Regional pronunciation variations were transcribed faithfully, with no corrections for mispronunciations. Additionally, numbers were transcribed as pronounced, maintaining consistency with the intended meaning of the speech.
@@ 3.2 Statistics
As shown in Table 3, our dataset consists of three subsets: training (317 speakers), validation (39 speakers), and test (41 speakers), with no overlap between speakers across the subsets. We further analyze the distribution of speakers based on age, gender, birthplace, accent, and recording device.
Figure 1 illustrates the age and gender distribution in our dataset. Due to the challenges in recruiting younger participants, the number of speakers decreases with younger age, while the gender distribution remains balanced across all age groups.
Figure 2 shows the distribution of utterance lengths and total speaking duration per speaker. Most utterances last between 1 and 5 seconds, with very few exceeding 10 seconds. The majority of speakers have a total speaking duration between 200 and 600 seconds, which is crucial for developing ASR systems tailored to young children.
Figure 3 shows the geographic distribution, covering 22 out of China’s 34 provincial-level administrative divisions. Although recruitment was challenging, we aimed for broad regional representation. Shanxi has the highest number of participants (136), followed by Jiangsu (40) and Henan (39). Provinces like Shaanxi, Shandong, and Hunan also contribute significantly. While some regions, such as Gansu, Heilongjiang, and Chongqing, have fewer participants, their inclusion highlights the dataset’s comprehensive geographic coverage.
Additionally, Figure 4 visualizes the distribution of speaker accents and recording devices. Accents are categorized into three levels: heavy (H), moderate (M), and light (L). The majority of speakers exhibit light accent variation, with only around 4% categorized as having moderate or heavy accents. We also ensured a balanced representation of iPhone and Android devices to support practical ASR system requirements.
@@
Wav2vec 2.0 (B) Enc Waveform 368M Self-sup. 10K
Wav2vec 2.0 (L) Enc Waveform 1,215M Self-sup. 10K
HuBERT (B) Enc Waveform 369M Self-sup. 10K
HuBERT (L) Enc Waveform 1,216M Self-sup. 10K
@@ 4.1 Speech recognition
For child speech recognition, we trained several baseline models from scratch and fine-tuned pre-trained models to evaluate performance on our dataset. For metrics, we employ Character Error Rate (CER, %), which is computed by the following equation:
$CER=\frac{S+D+I}{N},$ (1)
where S, D, I denote the numbers of substitutions, deletions and insertions, respectively. N represents the total number of characters in the reference. A system with a lower CER is generally considered superior in terms of character-level transcription accuracy.
@@ 4.1.1 Baselines trained from scratch
We utilize the open-source Wenet toolkit (Yao et al., 2021) to train ASR models from scratch. Three architectures are chosen: Transformer (Vaswani, 2017), Conformer (Gulati et al., 2020), and Paraformer (Gao et al., 2022). These models incorporate different approaches, including Connectionist Temporal Classification (CTC) (Graves et al., 2006), RNN-Transducer (RNN-T) (Graves, 2012), and attention based encoder-decoder (AED) (Chorowski et al., 2014; Chan et al., 2015).
@@
  • Conformer: The Conformer (Gulati et al., 2020) model integrates convolutions with self-attention for ASR, sandwiched between two feed-forward layers. For Conformer, we trained two models using both CTC and RNN-T loss functions respectively, following the Wenet recipe.
  • Paraformer: Proposed by Gao et al. (Gao et al., 2022), Paraformer is a fast and accurate parallel transformer model. It uses a continuous integrate-and-fire (CIF) (Dong and Xu, 2020) predictor to estimate the number of tokens and generate hidden representations.
@@ 4.1.4 Results of fine-tuning pre-trained models
Table 7 presents CER results for Conformer-WenetSpeech (CW) and Whisper models under zero-shot and fine-tuning settings. Fine-tuning results in substantial CER improvements for all supervised models. Despite Whisper’s large parameter size and extensive training data, the limited size of our dataset causes Whisper-medium to perform slightly worse than Whisper-Small after fine-tuning. Overall, CW achieves the best performance in both zero-shot and fine-tuned settings, highlighting its robust ASR capabilities learned from WenetSpeech.
@@ 4.2.2 Speaker verification baselines
In this study, three popular speaker embedding extractors, pre-trained on VoxCeleb (Nagrani et al., 2017), were fine-tuned on our dataset: x-vector [^5] https://huggingface.co/speechbrain/spkrec-xvect-voxceleb (Snyder et al., 2018), ECAPA-TDNN [^6] https://huggingface.co/speechbrain/spkrec-ecapa-voxceleb (Desplanques et al., 2020), and ResNet-TDNN [^7] https://huggingface.co/speechbrain/spkrec-resnet-voxceleb (Villalba et al., 2020). These models were implemented using the SpeechBrain (Ravanelli et al., 2021) toolkit and fine-tuned for 40 epochs. The embeddings extracted from the verification trials were then used to evaluate the models’ performance on the speaker verification task.
@@ 4.2.3 Results of speaker verification
Table 8 summarizes the performance of the models on the dataset, with both PLDA and Cosine Similarity evaluated using EER and minDCF metrics. Two key insights emerge from the results: First, the dataset proves to be well-suited for speaker-related tasks, as indicated by the strong performance of the three fine-tuned baseline models. However, the underdeveloped vocal characteristics of young children present challenges, potentially masking gender-related features and other distinguishing attributes. Second, due to the relatively small size of the dataset, the larger ECAPA-TDNN model underperformed compared to ResNet and x-vector, likely due to overfitting. Therefore, when applying this dataset to speaker verification tasks, particular attention should be given to enhancing the model’s generalization capability.
@@ 5 Conclusion
In conclusion, this paper introduces a valuable Mandarin speech dataset specifically designed for young children aged 3 to 5, addressing a crucial gap in ASR resources for this age group. Comprising 41.25 hours of speech data from 397 speakers across diverse provinces in China, the dataset ensures balanced gender representation and board geographic coverage. Our evaluations of ASR models and speaker verification show significant improvements, highlighting the dataset’s effectiveness in advancing children’s speech technology. This work represents a significant contribution to Mandarin child speech research and holds great promise for applications in educational technology and child-computer interaction. The dataset is freely available for academic use, supporting further advancements in the field.

### v3

@@ Abstract
Automatic speech recognition (ASR) systems have advanced significantly with models like Whisper, Conformer, and self-supervised frameworks such as Wav2vec 2.0 and HuBERT. However, developing robust ASR models for young children’s speech remains challenging due to differences in pronunciation, tone, and pace compared to adult speech. In this paper, we introduce a new Mandarin speech dataset focused on children aged 3 to 5, addressing the scarcity of resources in this area. The dataset comprises 41.25 hours of speech with carefully crafted manual transcriptions, collected from 397 speakers across various provinces in China, with balanced gender representation. We provide a comprehensive analysis of speaker demographics, speech duration distribution and geographic coverage. Additionally, we evaluate ASR performance on models trained from scratch, such as Conformer, as well as fine-tuned pre-trained models like HuBERT and Whisper, where fine-tuning demonstrates significant performance improvements. Furthermore, we assess speaker verification (SV) on our dataset, showing that, despite the challenges posed by the unique vocal characteristics of young children, the dataset effectively supports both ASR and SV tasks. This dataset is a valuable contribution to Mandarin child speech research. The dataset is now open-source and freely available for all academic purposes on https://github.com/flageval-baai/ChildMandarin.
@@
Jiaming Zhou1, Shiyao Wang1, Shiwan Zhao1, Jiabei He1, Haoqin Sun1, Hui Wang1, Cheng Liu1, Aobo Kong1, Yujie Guo1, Xi Yang2, Yequan Wang2, Yonghua Lin2, Yong Qin1thanks: Yong Qin is the corresponding author. 1 College of Computer Science, Nankai University, 2 Beijing Academy of Artificial Intelligence, Beijing, China, Correspondence: zhoujiaming@mail.nankai.edu.cn, qinyong@nankai.edu.cn
@@ 2.1 Child Speech Recognition Corpora in Mandarin Chinese
Gao et al. (2012) collected the CASS CHILD dataset, which contains 631 hours of speech from 23 children aged 1 to 4 years. However, only about 80 hours of this dataset are labeled with transcriptions, and, critically, the dataset is not publicly accessible. This restricts its use in ASR experiments and highlights the difficulty of obtaining child speech corpora in Mandarin.
@@ 3.1 Dataset details
The dataset consists of 41.25 hours of speech data with carefully crafted character-level manual transcriptions, collected from Mandarin-speaking children aged 3 to 5 years. The gender distribution is balanced across all age groups. To ensure geographic coverage, speakers were selected from different regions of China. A total of 397 speakers participated, representing 22 out of 34 provincial-level administrative divisions. Accents were classified into three categories: heavy (H), moderate (M), and light (L).
Our data collection occurred in a conversational context to promote natural interaction, with parents present throughout the sessions to provide emotional comfort and support for the children. The recording content was unrestricted, focusing on age-appropriate daily communication, ensuring that children engaged in familiar and non-stressful activities.
Prior to data collection, informed consent was obtained from the parents or legal guardians of all participants. The consent process included detailed explanations of the study’s purpose, procedures, and the intended use of the data for academic research. Parents were explicitly informed about their right to withdraw consent at any time without any repercussions.
@@
Character-level manual annotations were performed by professional transcribers, who meticulously adhered to the audio content, including stutters, disfluencies, and developmental speech patterns. Regional pronunciation variations were transcribed faithfully. Additionally, numbers were transcribed as pronounced, maintaining consistency with the intended meaning of the speech.
@@ 3.2 Statistics
As shown in Table 3, our dataset consists of three subsets: training (317 speakers), validation (39 speakers), and test (41 speakers), with no overlap between speakers across the subsets. We further analyze the distribution of speakers based on age, gender, birthplace, accent and recording device.
The age and gender distribution in the dataset, depicted in Figure 1, highlights a decrease in the number of speakers as age decreases, which reflects the challenges in recruiting younger participants. Despite this, the gender distribution remains balanced across all age groups.
The distribution of utterance lengths and total speaking duration per speaker is presented in Figure 2. Most utterances are between 1 and 5 seconds long, with very few exceeding 10 seconds. Additionally, the majority of speakers have a total speaking duration between 200 and 600 seconds, which is essential for developing ASR systems tailored to young children.
The geographic distribution of speakers, spanning 22 of China’s 34 provincial-level administrative divisions, is summarized in Figure 3. Despite recruitment challenges, broad regional representation was achieved, with Shanxi contributing the highest number of participants (136), followed by Jiangsu (40) and Henan (39). Provinces such as Shaanxi, Shandong, and Hunan also contribute significantly. Although some regions, including Gansu, Heilongjiang, and Chongqing, have fewer participants, their inclusion enhances the dataset’s comprehensive geographic coverage.
Speaker accents and recording devices are analyzed in Figure 4. Accents are categorized into three levels: heavy (H), moderate (M), and light (L), with the majority of speakers exhibiting light accent variation. Only around 4% of speakers are categorized as having moderate or heavy accents. Furthermore, a balanced representation of iPhone and Android devices was achieved to support diverse ASR system requirements.
@@
Wav2vec 2.0 (B) Enc Waveform 97M Self-sup. 10K
Wav2vec 2.0 (L) Enc Waveform 318M Self-sup. 10K
HuBERT (B) Enc Waveform 97M Self-sup. 10K
HuBERT (L) Enc Waveform 319M Self-sup. 10K
@@ 4.1 Speech recognition
For child speech recognition, we trained several baseline models from scratch and fine-tuned pre-trained models to assess performance on our dataset. We use the Character Error Rate (CER, %) as the evaluation metric. Refer to the Appendix A for the complete hyperparameter configurations.
@@ 4.1.1 Baselines trained from scratch
We utilize the open-source Wenet toolkit (Yao et al., 2021) to train ASR models from scratch. Three architectures are chosen: Transformer (Vaswani, 2017), Conformer (Gulati et al., 2020), and Paraformer (Gao et al., 2022). These models incorporate different approaches, including Connectionist Temporal Classification (CTC) (Graves et al., 2006), RNN-Transducer (RNN-T) (Graves, 2012), and Attention-based encoder-decoder (AED) (Chorowski et al., 2014; Chan et al., 2015).
@@
  • Conformer: The Conformer (Gulati et al., 2020) integrates convolutions with self-attention for ASR. We trained two models using both CTC and RNN-T loss functions respectively, following the Wenet recipe.
  • Paraformer: Proposed by Gao et al. (Gao et al., 2022), Paraformer is a fast and accurate parallel transformer model.
@@ 4.1.4 Results of fine-tuning pre-trained models
Table 7 presents the CER results for Conformer-WenetSpeech (CW) and Whisper models in zero-shot and fine-tuning settings. Fine-tuning results in substantial CER improvements for all supervised models. Despite Whisper’s large parameter size and extensive training data, the limited size of our dataset causes Whisper-medium to perform slightly worse than Whisper-Small after fine-tuning. Overall, CW achieves the best performance in both zero-shot and fine-tuned settings, highlighting its robust ASR capabilities learned from WenetSpeech.
Figure 5: CER (%) comparison of zero-shot and fine-tuning methods using CW model across different age-gender groups
Method Zero-shot Fine-tuning
Age_Gender S (%) D (%) I (%) S (%) D (%) I (%)
3_F 13.87 3.63 1.11 9.03 2.04 0.69
3_M 34.78 5.45 3.97 26.80 4.35 2.11
4_F 6.02 0.57 0.30 3.94 0.53 0.15
4_M 11.55 1.80 1.14 8.92 1.46 0.91
5_F 9.25 2.36 0.95 7.02 1.31 0.45
5_M 16.70 4.36 1.81 14.32 3.04 1.23
Table 8: In-depth comparison of different error types (S: Substitutions, D: Deletions, I: Insertions) between zero-shot and fine-tuning methods using CW model across different age-gender groups
@@
4.1.5 Performance Analysis
Figure 5 shows ASR performance across age and gender groups on the CW model. CER decreases with age, with 3-year-olds exhibiting higher error rates than 5-year-olds, reflecting greater variability in younger children’s speech. Fine-tuning significantly reduces CER across all age groups, demonstrating its effectiveness in adapting models to children’s speech.
Gender Trends: Male speakers consistently exhibit higher CER than female speakers of the same age. This disparity may arise from greater pitch and articulation variability in young male children.
Error Types: We further investigate error types in Tabel 8. Substitutions dominate error types, followed by deletions and insertions. Younger children, particularly 3-year-olds, exhibit higher substitution and deletion rates, reflecting challenges in speech recognition.
In summary, age and gender notably influence ASR performance, with younger and male speakers posing greater challenges. Fine-tuning mitigates these issues, highlighting the importance of targeted adaptation strategies. Detailed utterance analysis can be found in Appendix B.
@@ 4.2.2 Speaker verification baselines
In this study, three popular speaker embedding extractors, pre-trained on VoxCeleb (Nagrani et al., 2017), were fine-tuned on our dataset: x-vector [^5] https://huggingface.co/speechbrain/spkrec-xvect-voxceleb (Snyder et al., 2018), ECAPA-TDNN [^6] https://huggingface.co/speechbrain/spkrec-ecapa-voxceleb (Desplanques et al., 2020), and ResNet-TDNN [^7] https://huggingface.co/speechbrain/spkrec-resnet-voxceleb (Villalba et al., 2020). These models were implemented using the SpeechBrain (Ravanelli et al., 2021) toolkit and fine-tuned for 40 epochs. The embeddings extracted from the verification trials were then used to evaluate the models’ performance on the SV task. Refer to the Appendix A for the complete hyperparameter configurations.
@@ 4.2.3 Results of speaker verification
Table 9 summarizes the performance of the models on the dataset, with both PLDA and Cosine Similarity evaluated using EER and minDCF metrics. Two key insights emerge from the results: First, the dataset proves to be well-suited for speaker-related tasks, as indicated by the strong performance of the three fine-tuned baseline models. However, the underdeveloped vocal characteristics of young children present challenges, potentially masking gender-related features and other distinguishing attributes. Second, due to the relatively small size of the dataset, the larger ECAPA-TDNN model underperformed compared to ResNet and x-vector, likely due to overfitting.
@@ 5 Conclusion
In conclusion, this paper introduces a valuable Mandarin speech dataset specifically designed for young children aged 3 to 5, addressing a crucial gap in ASR resources for this age group. Comprising 41.25 hours of speech data from 397 speakers across diverse provinces in China, the dataset ensures balanced gender representation and board geographic coverage. Our evaluations of ASR models and speaker verification show significant improvements, highlighting the dataset’s effectiveness in advancing children’s speech technology. This work represents a significant contribution to Mandarin child speech research and holds great promise for applications in educational technology and child-computer interaction.
@@
Ethics Statement
This study adhered to strict ethical standards to safeguard the well-being and rights of participants. Recordings were conducted in a conversational context to encourage natural interaction, with parents present to provide emotional support. The content was unrestricted, focusing on age-appropriate, familiar communication to ensure a stress-free environment.
Informed consent was obtained from parents or legal guardians, who were fully briefed on the study’s purpose, procedures, and data use for academic research. They were informed of their right to withdraw at any time. Each child received a fair compensation of 150 RMB (about \$20 USD), carefully calibrated to avoid undue influence.
To protect privacy, all data was anonymized, removing personal identifiers and replacing them with coded labels. The dataset is securely stored, with access restricted to authorized researchers for academic purposes. The publicly available dataset will be licensed to prohibit commercial use and ensure compliance with ethical research practices.
While the study posed minimal risks, measures such as parental presence and familiar settings were implemented to ensure children’s psychological comfort. This dataset aims to advance automatic speech recognition (ASR) technologies for underrepresented groups like young children. However, we recognize the potential misuse of ASR technologies and have taken steps to mitigate risks by restricting dataset use to academic research and promoting ethical applications.
In summary, this study emphasizes informed consent, privacy protection, fair compensation, and ethical use of the data, ensuring respect for participants’ rights and well-being while contributing responsibly to the research community.
Appendix A Experimental configurations
Encoder Decoder Batch size LR Warmup Epochs
Transformer CTC+ATT 32 1.00E-03 2500 100
Conformer CTC+ATT 32 1.00E-03 2500 100
Conformer RNN-T+ATT 32 1.00E-03 2500 100
Paraformer Paraformer 16 2.00E-03 2500 100
Table A.1: Hyperparameters for training ASR models from scratch.
Model Batch size Learning rate Epochs
CW 16 4.00E-04 100
Wav2vec 2.0 (B) 10 3.00E-05 70
Wav2vec 2.0 (L) 10 1.00E-04 70
HuBERT (B) 10 5.00E-05 70
HuBERT (L) 10 5.00E-05 70
Whisper 16 1.00E-5 $\sim$ 1.00E-6 20
Table A.2: Hyperparameters for fine-tuning pre-trained ASR models.
Model Batch size LR Schedule Init LR Base LR Epochs
ECAPA-TDNN 128 Cyclic 5.00E-03 1.00E-08 40
ResNet-TDNN 128 Cyclic 5.00E-03 1.00E-08 40
X-vector 128 Linear 5.00E-03 1.00E-04 40
Table A.3: Hyperparameters for training speaker verification models.
This section provides detailed configurations and hyperparameters used for training and fine-tuning the ASR and speaker verification (SV) models discussed in the paper. All experiments were conducted using four GTX 3090 or GTX 4090 GPUs over several hours.
A.1 ASR model training from scratch
The hyperparameters for training ASR models from scratch are summarized in Table A.1. These models were trained using the Wenet toolkit with the configurations shown below.
A.2 ASR model fine-tuning
Table A.2 presents the fine-tuning hyperparameters for pre-trained ASR models, including Wav2vec 2.0, HuBERT, Whisper, and Conformer-WenetSpeech. Fine-tuning was performed using the training subset of our dataset.
A.3 Speaker verification (SV) model training
Table A.3 provides the training configurations for speaker verification models, including ECAPA-TDNN, ResNet-TDNN, and X-vector. These models were trained and evaluated on our dataset for speaker verification tasks.
Appendix B Analysis of fine-Tuning performance on specific utterances
As presented in Figure B.1, the fine-tuning process significantly improved the ASR model’s performance across various utterances, with a clear reduction in character error rate (CER). In general, fine-tuning allowed the model to adapt to specific child speech variations, addressing common issues such as phoneme substitutions and mispronunciations. Despite these improvements, some residual errors were still observed, particularly for more complex or longer utterances. Overall, the results demonstrate the effectiveness of fine-tuning for enhancing ASR performance on child speech, though further optimization is necessary to fully address all challenges.
Figure B.1: Performance comparison of zero-shot and fine-tuned models on specific utterances